Calculate the wavelengths of sounds in 14°C air, for each frequencies of 700Hz and 10,000Hz.
Answer #1

Velocity of sound at temeprature t degree celsius is given by

v=331+0.6t

At temperature 14 degree celsius is

v=(331+0.6\times14)\ m/s=339.4\ m/s

Hence, wavelength for frequency 700 Hz is

\lambda=\frac{Velocity}{Frequency}=\frac{339.4}{700}\ m=0.485\ m

For 1000 Hz frequency

\lambda=\frac{Velocity}{Frequency}=\frac{339.4}{10000}\ m=0.0339\ m